2. Four charges. q, 5.00 x10'C. q-3.00 x C. q-2.00 x10°C. -6.00 x10 C, are situated at the corners of a square of side 4.00m. Find the potential at the center of the square?


1
Expert's answer
2022-02-20T15:45:15-0500

Given:

"q_1=5.00*10^{-7}\:{\rm C},\quad q_2=-3.00*10^{-7}\:\rm C"

"q_3=-2.00*10^{-7}\:{\rm C},\quad q_4=6.00*10^{-7}\:\rm C"

"r=\frac{\sqrt{2}}{2}*4.00\:\rm m=2.83\: m"


The net potential at the center of square


"V=V_1+V_2+V_3+V_4""V=\frac{k}{r}(q_1+q_2+q_3+q_4)""V=\frac{9*10^9}{2.83}(5.00-3.00-2.00+6.00)*10^{-7}=1910\:\rm V"